OpenStudy (anonymous):

I need to use elimination to solve this problem idk what im doing wrong. my answer was (50,21) ...... not quite right i know.. {2x – 5y = -5 {-2x + 8y = -58

OpenStudy (welshfella):

if you add the 2 equations the terms in x are eliminated and you get 3y = -63 so y = -63/3

OpenStudy (welshfella):

so what is y?

OpenStudy (anonymous):

y=21

OpenStudy (welshfella):

no dont forget the negative!!

OpenStudy (anonymous):

oh y=-21

OpenStudy (welshfella):

yes now plug y = -21 into 2x - 5y = -5 to get value of x

OpenStudy (welshfella):

remember that - time - = +

OpenStudy (welshfella):

- times - = +

OpenStudy (anonymous):

2x-5(-21)=-5

OpenStudy (welshfella):

yea

OpenStudy (anonymous):

2x+105=-5

OpenStudy (anonymous):

2x=-110 right?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

now just divide by 2

OpenStudy (anonymous):

x=55

OpenStudy (welshfella):

No - NEGATIVE again!

OpenStudy (anonymous):

(55,-21) is the answer then

OpenStudy (anonymous):

oh dang it so (-55,-21)

OpenStudy (welshfella):

right

OpenStudy (anonymous):

thankyou

OpenStudy (welshfella):

yw